OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

3x-y = -7. we can do by 2 ways. one way is the two points should satisfy the equation. so we can get the line from options easily. other way is y-y1 = m(x-x1)=> y-1 = [(10-1)/ (1+2)] (x+2) => 3x-y = -7